OpenStudy (anonymous):

Factor completely 9x2 + 6x + 1 Answer None of the above (1 – 3x)2 (3x – 1)(3x + 1) (3x + 1)2

OpenStudy (anonymous):

When a quadratic expression is in the form \[ax^2+bx+c\], then you should try to find 2 numbers A, and B, such that: \[A*B=a*c\]\[A+B=b\]

jhonyy9 (jhonyy9):

(1-3x)(1-3x)=1-3x-3x+9x2=9x2 -6x +1

jhonyy9 (jhonyy9):

(3x+1)(3x+1)=9x2 +6x +1 so this is right sure (3x+1) squared